# MM-Ego: Towards Building Egocentric Multimodal LLMs

### v1

@@ Abstract
This research aims to comprehensively explore building a multimodal foundation model for egocentric video understanding. To achieve this goal, we work on three fronts. First, as there is a lack of QA data for egocentric video understanding, we develop a data engine that efficiently generates 7M high-quality QA samples for egocentric videos ranging from 30 seconds to one hour long, based on human-annotated data. This is currently the largest egocentric QA dataset. Second, we contribute a challenging egocentric QA benchmark with 629 videos and 7,026 questions to evaluate the models’ ability in recognizing and memorizing visual details across videos of varying lengths. We introduce a new de-biasing evaluation method to help mitigate the unavoidable language bias present in the models being evaluated. Third, we propose a specialized multimodal architecture featuring a novel “Memory Pointer Prompting” mechanism. This design includes a global glimpse step to gain an overarching understanding of the entire video and identify key visual information, followed by a fallback step that utilizes the key visual information to generate responses. This enables the model to more effectively comprehend extended video content. With the data, benchmark, and model, we successfully build MM-Ego, an egocentric multimodal LLM that shows powerful performance on egocentric video understanding. $\faApple\faApple{}^{\text{\faApple}}$ork done during an internship at Apple. First Authors. Senior Authors.
Figure 1: We introduce a foundation model for egocentric video understanding, contributing from three key perspectives: (a) a data engine that can automatically transform human narrations into 7 million egocentric QA samples, (b) a multimodal language model designed for egocentric video comprehension, and (c) the curation of a challenging egocentric video understanding benchmark.
@@ 1 Introduction
However, research on understanding egocentric videos remains in its early stages, with previous research primarily centered on specialized tasks such as story summarization (Lee et al., 2012), hand-object relationship understanding (Cai et al., 2016), action classification (Cartas et al., 2017; Li et al., 2021), and temporal or spatial grounding (Grauman et al., 2022). In contrast, works focusing on developing a more general egocentric video understanding model capable of complex understanding remain rare. Despite that video multimodal large language models (MLLMs) demonstrate strong video understanding and reasoning ability (Zhang et al., 2023; Wang et al., 2024; Lin et al., 2024; Zhang et al., 2024b), most of these works are unsuitable for egocentric video understanding from data, benchmark, and model design perspectives.
@@
$(a)$ From a data standpoint, although many MLLMs use some egocentric videos from ActivityNet (Yu et al., 2019), Ego4D (Grauman et al., 2022), and Charades (Sigurdsson et al., 2018) in their training, they have not been trained on large-scale egocentric video datasets, which inherently restricts their ability to comprehend lengthy first-person videos and accurately extract visual details. While Ego4D (Grauman et al., 2022) offers valuable human-annotated videos and labels for certain egocentric video understanding tasks, particularly episodic memory (which assesses a model’s ability to retain visual details in such videos), its annotations are not structured for generating language responses, making them unsuitable for training MLLMs. Therefore, a large-scale egocentric video QA corpus is still needed. $(b)$ In terms of benchmarking, exisiting video QA benchmarks either focus on shorter videos – such as EgoSchema (Mangalam et al., 2023) and QaEgo4D, which evaluate using around 3-minute and 8-minute videos, respectively – or concentrate on Internet video content (e.g., Video-MME (Fu et al., 2024)). This creates a notable gap in egocentric video understanding benchmarks that encompass videos ranging from seconds to an hour in length. $(c)$ From a model design perspective, previous video MLLMs have primarily addressed long videos in two ways. The first approach involves uniformly sampling a limited number of video frames as visual input, as seen in Li et al. (2024); Lin et al. (2024). Despite its simplicity, this approach achieves better performance among open-source models on public video benchmarks (Fu et al., 2024), largely because its design ensures high training efficiency and good scaling properties. The second approach involves feeding a large volume of visual tokens into the transformer backbone and employing engineering techniques, such as tensor parallelism and sequence parallelism (Xue et al., 2024; Zhang et al., 2024a), to facilitate training with millions of visual tokens in context. However, these long-context transformers suffer from slow training speeds and small overall batch sizes, which hinder performance improvements given the constraints of computational resources and training time. Intuitively, even humans cannot remember every detail of an hour-long video. We believe a more effective approach is to understand the video progressively: first get an overview of the entire video, then focus on specific details with particular questions in mind.
@@
($i$) Data. To scale training data for MLLMs with egocentric understanding ability, we develop an efficient data engine, using a “narration to egocentric QA” strategy, to automatically synthesize a large-scale egocentric QA dataset based on video narration data. Notably, rather than relying on existing vision-language models as labelers, we generate egocentric QAs based on the human-annotated fine-grained video clip narrations. This approach ensures that our data quality is not constrained by the limitations of existing vision-language labeling models. In this way, we create the first large-scale egocentric QA dataset, consisting of over 7 million egocentric QA samples that span video lengths from seconds to over an hour. This dataset enables the training of models to recognize and retain visual details from egocentric videos.
@@ 2 Method
The workflow of the data engine is illustrated in Figure 2. By organizing sequential video clips {Clip 1, Clip 2, …, Clip N} and their corresponding narrations {Narration 1, Narration 2, …, Narration N} in proper chronological order, we create comprehensive narration paragraphs that describe entire video sequences. We then employ a powerful text-only language model, i.e., GPT4-o, to generate diverse and confident QA pairs related to episodic memory based on these narration paragraphs. The language model is instructed to attach the index of the narration sentence upon which each QA pair is based. This indexing allows us to map each QA pair back to the corresponding time frames in the original videos, enabling the extraction of key frame information crucial for subsequent model training.
Applying this data engine to the extensive Ego4D dataset allows us to efficiently scale the creation of egocentric QA data. We partition the dataset into training and testing sets according to the official Ego4D episodic memory task. The egocentric QA dataset provides more than 7 million QA samples in 938K multi-turn conversations. The data encompasses videos of varying durations, ranging from 30 seconds to 1 hour, as illustrated in Figure 3. To ensure comprehensive coverage and prevent bias towards shorter videos, we balance the number of conversations across different video lengths in training. This is the first large-scale egocentric QA dataset featuring videos of such extended ranges of duration.
@@ 2.2 MM-Ego Model
Our modeling goal is to develop an MLLM for handling egocentric videos, which are lengthy and rich in visual details. On the one hand, frame-level information is necessary to capture the full content of the video, as skipping frames during sampling could result in a significant loss of visual details. On the other hand, processing all visual tokens generated by the visual encoder is computationally challenging for the transformer model. For instance, if each image is encoded into 729 visual embeddings (tokens), the total number of visual embeddings for a 300-frame video would be 218,700. However, most MLLMs are trained with a context length of less than 10,000 tokens (Li et al., 2024). Taking these factors into account, we introduce the MM-Ego model, which is built for handling a large volume of egocentric video frames while maintaining manageable computational costs within the transformer backbone. MM-Ego introduces an innovative Memory Pointer Prompting mechanism, which operates in two main steps: global glimpse and fallback. We will introduce the details of MM-Ego in the following sections.
@@ 2.2.1 Visual and Textual Embedding
Given an input video and the question, the first step is to embed them into visual and textual embeddings separately for later processing. We begin by uniformly sampling the video into up to $N$ frames, where $N$ can be in the range of hundreds. Then, we extract per-frame visual feature maps from these frames using a robust vision encoder, SigLIP-so400m (Zhai et al., 2023). Following the method outlined by Li et al. (2024), we apply a 2-layer MLP to project the visual feature maps to the LLM embedding space and use average pooling to reduce the height and width of the visual feature maps by a factor of two and flatten the height and width dimension, resulting in $N$ relatively high-resolution visual embeddings $\textbf{V}^{i}\in\mathbb{R}^{T\times C}$, $i\in[1,N]$} where $T$ is the embedding length and $C$ is the embedding dimension. For the textual embedding, since we use Qwen2 (Yang et al., 2024) as the LLM, we use its tokenizer and embedding layer to transform the input text into textual embeddings. For question $q$, we denote the corresponding textual question embedding as $\textbf{E}_{\text{que}}^{q}\in\mathbb{R}^{T_{q}\times C}$, $q\in[1,Q]$} where $Q$ is the total number of questions and $T_{q}$ is the embedding length of question $q$.
@@ 3.1 EgoMemoria Benchmark
To evaluate the performance of egocentric MLLMs, especially in terms of episodic memory ability, we propose a new benchmark called EgoMemoria. Specifically, we generate memory-related questions and answers from human-annotated narrations in the validation set of the Ego4D dataset. To ensure diversity, for each video we only generate a limited number of questions. We divide the videos into seven different length ranges: 0.5 to 1 min, 1 to 2 min, 2 to 4 min, 4 to 10 min, 10 to 20 min, 20 to 40 min, and 40 to 60 min. We aim to balance the number of samples in different video lengths. The distribution of videos and corresponding question-answer pairs (QAs) for each category is shown in Table 2. Furthermore, we group these video lengths into three broader categories: short (0.5 to 2 min), medium (2 to 20 min), and long (20 to 60 min). In total, we collect 629 videos with 7,026 questions. The most frequently occuring verbs and nouns in the questions are visualized in Figure 5.
@@ 3.2 Experimental Setup
Implementation Details. The model is trained for one epoch with a base learning rate of $1\times 10^{-5}$, using a cosine scheduler. The batch size is set to 128. We sample a maximum of 300 frames ($N=300$) and select 32 visual embeddings in the proposed memory pointer prompting mechanism. By default, we set the explore-exploit balancing parameter $\alpha$ to $0.1$.
Pretrained Models. Our MM-Ego model is initialized from LLaVA-OV 7B (Li et al., 2024), a state-of-the-art MLLM known for its good performance on general multimodal understanding tasks. Following the same architecture, we use the SigLip-so400M ViT (Zhai et al., 2023) as the visual encoder for embedding video frames and Qwen2-7B (Yang et al., 2024) as the LLM architecture.
@@ 3.3 Main Results
We first conduct experiments on our EgoMemoria benchmark, primarily comparing three models: LLaVA-OV (Li et al., 2024), its fine-tuned version using our MM-Ego SFT data mixture (referred to as “Ego SFT”), and our MM-Ego model, which incorporates the proposed Memory Pointer Prompting mentioned in Section 2.2.2. We show the EgoMemoria accuracy in the first row of Table 3. We observe a significant improvement in the model’s performance on egocentric QAs after training on our MM-Ego data mixture, attributed to the rich egocentric knowledge provided by our curated egocentric QA training data. Moreover, leveraging the MM-Ego model architecture further enhances performance, thanks to the effective Memory Pointer Prompting mechanism.
@@
To better understand the capability of MM-Ego, we compare its performance with state-of-the-art video MLLMs on EgoMemoria and prevalent large-scale video QA benchmarks, including the long egocentric video understanding benchmark EgoSchema (Mangalam et al., 2023) and the Internet-video-based long-video understanding benchmark Video-MME (Fu et al., 2024). The results are shown in Table 4. On EgoMemoria, GPT-4o is evaluated using 32 uniformly sampled frames from the videos, while other models follow their respective official inference settings. The MDA on EgoMemoria is computed using the debiased subsets used in Table 3. Notably, MM-Ego exhibits the highest performance on EgoMemoria, particularly in the ‘Medium’ and ‘Long’ classes. On the EgoSchema benchmark, our model achieves a substantial performance gain of +8.18 over the previous state-of-the-art open-source model, underscoring the effectiveness of both our data and model design for egocentric understanding. Additionally, on the challenging Internet video understanding Video-MME benchmark, our model is on par with the leading model of similar parameter size. These results showcase MM-Ego’s capability in egocentric video understanding while preserving its general video comprehension abilities.
@@ 4 Related Work
Recent advancements in Large Language Models (OpenAI, 2023; Touvron et al., 2023) have sparked significant interest in developing Multimodal Large Language Models (MLLMs) that combine the language understanding capabilities of LLMs with multi-modal perception abilities (Alayrac et al., 2022; Dai et al., 2023; Zhu et al., 2023; McKinzie et al., 2024). For video-based MLLMs, most works follow a structure akin to image-based MLLMs. To handle the large volume of video frames, some methods reduce the number of frames (Zhang et al., 2023; Wang et al., 2024; Maaz et al., 2024; Xu et al., 2024), which results in the loss of many visual details. Others extend the LLMs’ context length by employing parallel techniques (Xue et al., 2024), but this often leads to low training efficiency. Unlike these approaches, our method preserves global awareness of the entire video, allows for attention to visual details, and is efficiently trainable.
@@
While the growing field of egocentric video understanding is still in its infancy, there have been many influential works. For a comprehensive overview of egocentric vision please refer to Plizzari et al. (2024). On the data/benchmark side, representative works include Ego4D (Grauman et al., 2022), Ego-Exo4D (Grauman et al., 2024), and EPIC-KITCHENS-100 (Damen et al., 2018). When also considering language, prior work on egocentric video-language benchmarks include QaEgo4D (Bärmann & Waibel, 2022) and EgoSchema (Mangalam et al., 2023). For understanding long egocentric videos, prior modeling efforts include GroundVQA (Di & Xie, 2024), Encode-Store-Retrieve (Shen et al., 2023), and R-VLM (Xu et al., 2023). However, most previous works focus on classic video understanding tasks such as activity recognition and temporal grounding, and hence they do not involve a language model. In contrast, we propose to develop a multimodal LLM to tackle open-ended egocentric video-language understanding.
5 Discussion
Limitation and Future Work. While MM-Ego demonstrates a strong ability in egocentric understanding, there is still room for further improvement. On the data and benchmark side, we can introduce more diverse egocentric understanding corpus. For the model itself, we plan to enhance its capacity to process a larger number of frames, such as at the order of thousands, to better handle longer or even always-on egocentric videos.
Conclusion. In this paper, we make three key contributions towards the development of egocentric foundation model: the creation of the first large-scale egocentric QA training dataset, the introduction of a novel model designed for effective long egocentric video comprehension, and the establishment of the EgoMemoria benchmark for assessing models’ ability to capture visual details from egocentric videos. We hope that these efforts will benefit further research on egocentric MLLMs.

### v2

@@ Abstract
This research aims to comprehensively explore building a multimodal foundation model for egocentric video understanding. To achieve this goal, we work on three fronts. First, as there is a lack of QA data for egocentric video understanding, we automatically generate 7M high-quality QA samples for egocentric videos ranging from 30 seconds to one hour long in Ego4D (Grauman et al., 2022) based on human-annotated data. This is one of the largest egocentric QA datasets. Second, we contribute a challenging egocentric QA benchmark with 629 videos and 7,026 questions to evaluate the models’ ability in recognizing and memorizing visual details across videos of varying lengths. We introduce a new de-biasing evaluation method to help mitigate the unavoidable language bias present in the models being evaluated. Third, we propose a specialized multimodal architecture featuring a novel “Memory Pointer Prompting” mechanism. This design includes a global glimpse step to gain an overarching understanding of the entire video and identify key visual information, followed by a fallback step that utilizes the key visual information to generate responses. This enables the model to more effectively comprehend extended video content. With the data, benchmark, and model, we build MM-Ego, an egocentric multimodal LLM that shows powerful performance on egocentric video understanding. $\faApple\faApple{}^{\text{\faApple}}$ork done during an internship at Apple. First Authors. Senior Authors.
Figure 1: We introduce a foundation model for egocentric video understanding, contributing from three key perspectives: (a) 7 million egocentric QA samples generated from human narrations via a data engine, (b) a multimodal language model designed for egocentric video comprehension, and (c) the curation of a challenging egocentric video understanding benchmark.
@@ 1 Introduction
However, research on understanding egocentric videos remains in its early stages, with previous research primarily centered on specialized tasks such as story summarization (Lee et al., 2012), hand-object relationship understanding (Cai et al., 2016), action classification (Cartas et al., 2017; Li et al., 2021), and temporal or spatial grounding (Grauman et al., 2022). In contrast, works focusing on developing a more general egocentric video understanding model capable of complex understanding remain rare. Despite that video multimodal large language models (MLLMs) demonstrate strong video understanding and reasoning ability (Zhang et al., 2023a; Wang et al., 2024b; Lin et al., 2024; Zhang et al., 2024b), most of these works are unsuitable for egocentric video understanding from data, benchmark, and model design perspectives.
@@
$(a)$ From a data standpoint, although many MLLMs use some egocentric videos from ActivityNet (Yu et al., 2019), Ego4D (Grauman et al., 2022), and Charades (Sigurdsson et al., 2018) in their training, they have not been trained on large-scale egocentric video datasets, which inherently restricts their ability to comprehend lengthy first-person videos and accurately extract visual details. While Ego4D (Grauman et al., 2022) offers valuable human-annotated videos and labels for certain egocentric video understanding tasks, particularly episodic memory (which assesses a model’s ability to retain visual details in such videos), its annotations are not structured for generating language responses, making them unsuitable for training MLLMs. Therefore, a large-scale egocentric video QA corpus is still needed. $(b)$ In terms of benchmarking, exisiting video QA benchmarks either focus on shorter videos – such as EgoSchema (Mangalam et al., 2023) and QaEgo4D, which evaluate using around 3-minute and 8-minute videos, respectively – or concentrate on Internet video content (e.g., Video-MME (Fu et al., 2024)). This creates a notable gap in egocentric video understanding benchmarks that encompass videos ranging from seconds to an hour in length. $(c)$ From a model design perspective, previous video MLLMs have primarily addressed long videos in two ways. The first approach involves uniformly sampling a limited number of video frames as visual input, as seen in Li et al. (2024a); Lin et al. (2024). Despite its simplicity, this approach achieves better performance among open-source models on public video benchmarks (Fu et al., 2024), largely because its design ensures high training efficiency and good scaling properties. The second approach involves feeding a large volume of visual tokens into the transformer backbone and employing engineering techniques, such as tensor parallelism and sequence parallelism (Xue et al., 2024; Zhang et al., 2024a), to facilitate training with millions of visual tokens in context. However, these long-context transformers suffer from slow training speeds and small overall batch sizes, which hinder performance improvements given the constraints of computational resources and training time. Intuitively, even humans cannot remember every detail of an hour-long video. We believe a more effective approach is to understand the video progressively: first get an overview of the entire video, then focus on specific details with particular questions in mind.
@@
($i$) Data. To scale training data for MLLMs with egocentric understanding ability, we develop an efficient data engine, using a “narration to egocentric QA” strategy, to automatically synthesize a large-scale egocentric QA dataset based on video narration data. Notably, rather than relying on existing vision-language models (VLMs) as labelers, we generate egocentric QAs based on the human-annotated fine-grained video clip narrations. This approach, conceptually related to (Di & Xie, 2024; Li et al., 2024b), ensures that our data quality is not constrained by the limitations of existing labeling VLMs. In this way, we create one of the first large-scale egocentric QA datasets, consisting of over 7 million egocentric QA samples that span video lengths from seconds to over an hour. This dataset enables the training of models to recognize and retain visual details from egocentric videos.
@@ 2 Method
The workflow of the data engine is illustrated in Figure 2. By organizing sequential video clips {Clip 1, Clip 2, …, Clip N} and their corresponding narrations {Narration 1, Narration 2, …, Narration N} in proper chronological order, we create comprehensive narration paragraphs that describe entire video sequences. We then employ a powerful text-only language model, i.e., GPT-4o, to generate diverse and confident QA pairs related to episodic memory based on these narration paragraphs. The language model is instructed to attach the index of the narration sentence upon which each QA pair is based. This indexing allows us to map each QA pair back to the corresponding time frames in the original videos, enabling the extraction of key frame information crucial for subsequent model training.
Applying this data engine to the extensive Ego4D dataset allows us to efficiently scale the creation of egocentric QA data. We partition the dataset into training and testing sets according to the official Ego4D episodic memory task. The egocentric QA dataset provides more than 7 million QA samples in 938K multi-turn conversations. The data encompasses videos of varying durations, ranging from 30 seconds to 1 hour, as illustrated in Figure 3. To ensure comprehensive coverage and prevent bias towards shorter videos, we balance the number of conversations across different video lengths in training. This is one of the first large-scale egocentric QA datasets featuring videos of such extended ranges of duration.
@@ 2.2 MM-Ego Model
Our modeling goal is to develop an MLLM for handling egocentric videos, which are lengthy and rich in visual details. On the one hand, frame-level information is necessary to capture the full content of the video, as skipping frames during sampling could result in a significant loss of visual details. On the other hand, processing all visual tokens generated by the visual encoder is computationally challenging for the transformer model. For instance, if each image is encoded into 729 visual embeddings (tokens), the total number of visual embeddings for a 300-frame video would be 218,700. However, most MLLMs are trained with a context length of less than 10,000 tokens (Li et al., 2024a). Taking these factors into account, we introduce the MM-Ego model, which is built for handling a large volume of egocentric video frames while maintaining manageable computational costs within the transformer backbone. MM-Ego introduces an innovative Memory Pointer Prompting mechanism, which operates in two main steps: global glimpse and fallback. We will introduce the details of MM-Ego in the following sections.
@@ 2.2.1 Visual and Textual Embedding
Given an input video and the question, the first step is to embed them into visual and textual embeddings separately for later processing. We begin by uniformly sampling the video into up to $N$ frames, where $N$ can be in the range of hundreds. Then, we extract per-frame visual feature maps from these frames using a robust vision encoder, SigLIP-so400m (Zhai et al., 2023). Following the method outlined by Li et al. (2024a), we apply a 2-layer MLP to project the visual feature maps to the LLM embedding space and use average pooling to reduce the height and width of the visual feature maps by a factor of two and flatten the height and width dimension, resulting in $N$ relatively high-resolution visual embeddings $\textbf{V}^{i}\in\mathbb{R}^{T\times C}$, $i\in[1,N]$} where $T$ is the embedding length and $C$ is the embedding dimension. For the textual embedding, since we use Qwen2 (Yang et al., 2024) as the LLM, we use its tokenizer and embedding layer to transform the input text into textual embeddings. For question $q$, we denote the corresponding textual question embedding as $\textbf{E}_{\text{que}}^{q}\in\mathbb{R}^{T_{q}\times C}$, $q\in[1,Q]$} where $Q$ is the total number of questions and $T_{q}$ is the embedding length of question $q$.
@@ 3.1 EgoMemoria Benchmark
To evaluate the performance of egocentric MLLMs, especially in terms of episodic memory ability, we propose a new benchmark called EgoMemoria. Specifically, we generate memory-related questions and answers from human-annotated narrations in the validation set of the Ego4D dataset. To ensure diversity, for each video we only generate a limited number of questions. We divide the videos into seven different length ranges: 0.5 to 1 min, 1 to 2 min, 2 to 4 min, 4 to 10 min, 10 to 20 min, 20 to 40 min, and 40 to 60 min. We aim to balance the number of samples in different video lengths. The distribution of videos and corresponding question-answer pairs (QAs) for each category is shown in Table 2. Furthermore, we group these video lengths into three broader categories: short (0.5 to 2 min), medium (2 to 20 min), and long (20 to 60 min). In total, we collect 629 videos with 7,026 questions. The most frequently occurring verbs and nouns in the questions are visualized in Figure 5.
@@ 3.2 Experimental Setup
Implementation Details. The model is trained for 1 epoch with a base learning rate of $1\times 10^{-5}$, using a cosine scheduler. The batch size is set to 128. We sample a maximum of 300 frames ($N=300$) and select 32 visual embeddings in the proposed memory pointer prompting mechanism. By default, we set the explore-exploit balancing parameter $\alpha$ to $0.1$. Greedy decoding is used in generation.
Pretrained Models. Our MM-Ego model is initialized from LLaVA-OV 7B (Li et al., 2024a), a state-of-the-art MLLM known for its good performance on general multimodal understanding tasks. Following the same architecture, we use the SigLip-so400M ViT (Zhai et al., 2023) as the visual encoder for embedding video frames and Qwen2-7B (Yang et al., 2024) as the LLM architecture.
@@ 3.3 Main Results
We first conduct experiments on our EgoMemoria benchmark, primarily comparing three models: LLaVA-OV (Li et al., 2024a), its fine-tuned version using our MM-Ego SFT data mixture (referred to as “Ego SFT”), and our MM-Ego model, which incorporates the proposed Memory Pointer Prompting mentioned in Section 2.2.2. We show the EgoMemoria accuracy in the first row of Table 3. We observe a significant improvement in the model’s performance on egocentric QAs after training on our MM-Ego data mixture, attributed to the rich egocentric knowledge provided by our curated egocentric QA training data. Moreover, leveraging the MM-Ego model architecture further enhances performance, thanks to the effective Memory Pointer Prompting mechanism.
@@
To better understand the capability of MM-Ego, we compare its performance with state-of-the-art video MLLMs on EgoMemoria and prevalent large-scale video QA benchmarks, including the long egocentric video understanding benchmark EgoSchema (Mangalam et al., 2023) and the Internet-video-based long-video understanding benchmark Video-MME (Fu et al., 2024). The results are shown in Table 4. On EgoMemoria, GPT-4o is evaluated using 32 uniformly sampled frames from the videos, while other models follow their respective official inference settings. The MDA on EgoMemoria is computed using the debiased subsets used in Table 3. Notably, MM-Ego exhibits the highest performance on EgoMemoria, particularly in the ‘Medium’ and ‘Long’ classes. On the EgoSchema benchmark, our model achieves a substantial performance gain of +8.18 over the previous state-of-the-art open-source model (“LLaVA-NeXT-Video-32B-Qwen”), underscoring the effectiveness of both our data and model design for egocentric understanding. Additionally, on the challenging Internet video understanding Video-MME benchmark, our model is on par with the leading model of similar parameter size although our data mixture is less diverse compared with (Li et al., 2024a). These results showcase MM-Ego’s capability in egocentric video understanding while preserving its general video comprehension abilities.
@@ 4 Related Work
Recent advancements in Large Language Models (OpenAI, 2023; Touvron et al., 2023) have sparked significant interest in developing Multimodal Large Language Models (MLLMs) that combine the language understanding capabilities of LLMs with multi-modal perception abilities (Alayrac et al., 2022; Dai et al., 2023; Zhu et al., 2023; McKinzie et al., 2024). For video-based MLLMs, most works follow a structure akin to image-based MLLMs. To handle the large volume of video frames, some methods reduce the number of frames (Zhang et al., 2023a; Wang et al., 2024b; Maaz et al., 2024; Xu et al., 2024), which results in the loss of many visual details. Others extend the LLMs’ context length by employing parallel techniques (Xue et al., 2024), but this often leads to low training efficiency. Unlike these approaches, our method preserves global awareness of the entire video, allows for attention to visual details, and is efficiently trainable.
@@
While the growing field of egocentric video understanding is still in its infancy, there have been many influential works. For a comprehensive overview of egocentric vision please refer to Plizzari et al. (2024). On the data/benchmark side, representative works include Ego4D (Grauman et al., 2022), Ego-Exo4D (Grauman et al., 2024), and EPIC-KITCHENS-100 (Damen et al., 2018). When also considering language, prior work on egocentric video-language benchmarks include QaEgo4D (Bärmann & Waibel, 2022) and EgoSchema (Mangalam et al., 2023). For understanding long egocentric videos, prior modeling efforts include GroundVQA (Di & Xie, 2024), Encode-Store-Retrieve (Shen et al., 2023), and R-VLM (Xu et al., 2023). However, most previous works focus on classic video understanding tasks such as activity recognition and temporal grounding, and hence they do not involve a large language model for complex understanding. In contrast, we propose to develop an MLLM to tackle comprehensive egocentric video understanding.
5 Conclusion
In this paper, we make three key contributions towards the development of egocentric foundation models: the creation of a large-scale egocentric QA training dataset, the introduction of a novel model designed for effective long egocentric video comprehension, and the establishment of the EgoMemoria benchmark for assessing models’ ability to capture visual details from egocentric videos. We hope that these efforts will benefit further research on egocentric MLLMs.